Session: Feb 2004 (session 1)
Wavelength: 30cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Thu Feb 26 11:06:08 2004) Minor failures : TSYS
As planned, we observed the two linear polarisations. Additionally, no system temperature measurements were done (we had problems with the diode). No other known problems.

Urumqi (zhanghua at Thu Feb 26 11:25:54 2004) Minor failures : HIGH
Tsys was high for 184.49MHz.vc3 and vc4 LSB power exceed IF power.

Onsala (Michael Lindqvist at Fri Feb 27 10:36:17 2004) Success :
No known problems

Noto (P. Cassaro at Sun Feb 29 23:57:32 2004) Success :
No known problems